OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 7, 21, 63, 189, ...

OpenStudy (anonymous):

an = 7 • 3n an = 7 • 3n - 1 an = 3 • 7n - 1 an = 7 • 3n + 1

hartnn (hartnn):

is the first choice \(a_n=7 \times 3^n\) ??

OpenStudy (anonymous):

yeah i guess. 7 x 3n is the same thing

hartnn (hartnn):

3n and 3^n are not the same thing...

OpenStudy (anonymous):

well then is is actually 3n the way it is written

OpenStudy (anonymous):

no it is 3^n

hartnn (hartnn):

yesss.

hartnn (hartnn):

so, do you notice, in that sequence that if you multiply each term by 3, you get the next term ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so would it be 7 x 3^n +1

hartnn (hartnn):

lets check! for n=1 7x 3^(1+1) = 7*3^2 does not equal 7 because n=1 implies first term which should be 7. so, its not 7. 3^(n+1)

OpenStudy (anonymous):

so then it has to be 7 x 3^n

hartnn (hartnn):

lets check again! for n=1, 7 x 3^1 =.... ?

OpenStudy (anonymous):

21

hartnn (hartnn):

n=1 means first term, it should come out to be 7. does it ?

OpenStudy (anonymous):

yea so would it =147

OpenStudy (anonymous):

???????????????

hartnn (hartnn):

where does 147 come from ?? :O

OpenStudy (anonymous):

idk know just forget it. so the first term is 7, so would A be the answer

hartnn (hartnn):

for which of the 4 : an = 7 • 3n an = 7 • 3n - 1 an = 3 • 7n - 1 an = 7 • 3n + 1 is the 1st term =7 ?? you can find out just by plugging, n=1.

OpenStudy (anonymous):

an= 7x 3n

hartnn (hartnn):

a1 = 7x 3^1 = 21 does not =7 so its not 7x3n

OpenStudy (anonymous):

7 x 3n-1

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

thnx

hartnn (hartnn):

welcome :)